OpenStudy (anonymous):

What is the mean of the following set of numbers? 6, 1, 3, 6, 11, 14, 7, 10 11 7.25 6.5 3

OpenStudy (anonymous):

The mean is the average.

OpenStudy (anonymous):

If you mean every number on there then the mean is 6.2. If you mean just the top row of numbers then it is 7.25.

OpenStudy (anonymous):

58/8=7.25

OpenStudy (anonymous):

What is the median of the following set of numbers? 62, 66, 66, 69, 73, 75, 77, 77, 80, 80, 80, 84 66 75 77 76 80

OpenStudy (anonymous):

I didn't get this one

OpenStudy (anonymous):

The median is post to be the number in the middle but there isnt a middle

OpenStudy (anonymous):

Find the mean of the two middle numbers.

OpenStudy (anonymous):

75+77=152 152/2=76

OpenStudy (anonymous):

The median is 76. If there isn't a number up there that you can actually look at then you have to go even farther into the problem. P.S. Can we get medals for this? ;)